# Structure As Search: Unsupervised Permutation Learning for Combinatorial Optimization

### v1

@@ 6.1 Main Theorem
Consider the orbit of vertex 00 under this mapping:
@@
Sufficiency ($\Leftarrow$): Suppose $\gcd(k,n)=1$. Then by the argument above, the orbit of vertex 00 has size $\frac{n}{1}=n$. This means the sequence $\{0,k,2k,\ldots,(n-1)k\}$ modulo $n$ contains all distinct elements of $\{0,1,\ldots,n-1\}$.
@@
Figure 5: TSP-20: Mean vs. Coprime Shifts
Figure 6: TSP-50: Mean vs. Coprime Shifts
Figure 7: TSP-100: Mean vs. Coprime Shifts
Figure 8: Best mean tour length on TSP instances of different sizes as the number of coprime shifts increases. Using more shift combinations significantly reduces tour length, outperforming the Greedy Multi-Start baseline. Both Christofides and Beam Search bounds are shown for comparison.
@@ 6.2 Ensemble Training and Inference
Table 2: Detailed performance comparison of learning-based TSP solvers across different instance sizes (TSP-20/50/100). Metrics include average tour length and optimality gap (%). Results for baseline methods are taken from (Joshi et al., 2019). While all methods are evaluated on uniformly distributed TSP instances generated in a consistent manner, the test sets differ slightly across works. Consequently, lower average tour length does not necessarily imply a smaller optimality gap when comparing across methods. For example, our ensemble approach reports a slightly lower tour length than PtrNet Bello et al. (2016) on TSP-100 (8.30 vs. 8.31), yet achieves a smaller gap (7.10% vs. 7.03%) due to test set variation. This table is intended to provide a general impression of the performance gap range across representative models. Note that more sophisticated combinations of RL/SL with search heuristics exist, which may yield improved results beyond those reported here.
@@
PtrNet Vinyals et al. (2015) SL, G 3.88 1.15% 7.66 34.48% - -
PtrNet Bello et al. (2016) RL, G 3.89 1.42% 5.95 4.46% 8.30 6.90%
S2V Khalil et al. (2017) RL, G 3.89 1.42% 5.99 5.16% 8.31 7.03%
GAT Deudon et al. (2018) RL, G 3.86 0.66% 5.92 3.98% 8.42 8.41%
GAT Deudon et al. (2018) RL, G, 2-OPT 3.85 0.42% 5.85 2.77% 8.17 5.21%
GAT Kool et al. (2018) RL, G 3.85 0.34% 5.80 1.76% 8.12 4.53%
GCN Joshi et al. (2019) SL, G 3.86 0.60% 5.87 3.10% 8.41 8.38%
@@
Hamiltonian cycle ensemble UL, NAR 3.97 3.52% 5.93 4.41% 8.30 7.10%
@@
Figures 8, 8, and 8 demonstrate the effectiveness of this ensemble approach across 20, 50, and 100-node problems respectively. As the number of coprime shifts increases, the mean tour length decreases substantially, with dramatic improvements observed initially that gradually plateau. For TSP-20, using all $\varphi(20)=8$ coprime shifts reduces mean tour length from 4.06 to 3.97, surpassing both multi-start greedy and beam search performance. Similar trends are observed for TSP-50 and TSP-100, where the ensemble approach achieves mean tour lengths of 5.93 and 8.30 respectively when using all available coprime shifts. This demonstrates that leveraging multiple Hamiltonian cycle structures effectively mitigates the long tail problem while consistently improving solution quality. Here, each permutation learner in our framework is trained with respect to a fixed initial Hamiltonian cycle $\mathbb{V}^{k}$, which serves as a structural prior guiding solution formation. This initialization anchors the training process, focusing learning on the permutation of the initial Hamiltonian cycle $\mathbb{V}^{k}$, effectively biasing the model. Since different initial cycles encode distinct structural priors, using an ensemble of models with $\mathbb{V}^{k}$ promotes diversity and improves overall solution quality.
@@
Despite not using supervision or autoregressive decoding, our method achieves competitive results across all TSP sizes. Our Hamiltonian cycle ensemble approach significantly outperforms classical baselines, achieving optimality gaps of 3.52%, 4.41%, and 7.10% on TSP-20, TSP-50, and TSP-100 respectively, compared to greedy NN (all start cities) search’s 5.74%, 11.6%, and 14.8%. We also improve upon beam search variants as the problem size grows, with beam search achieving gaps of 3.13–3.92% on TSP-20, 17.0–17.9% on TSP-50, and 25.3–26.1% on TSP-100.
@@
Among learning-based methods, our approach demonstrates competitive performance with early models. We achieve comparable results to Pointer Networks and S2V. Our method also performs competitively with supervised method Joshi et al. (2019), achieving 7.10% optimality gap versus 8.38% on TSP-100. However, we do not yet match the best-performing models such as the attention model by Kool et al. Kool et al. (2018) or GAT with 2-OPT refinement Deudon et al. (2018), which benefit from reinforcement learning training and post-hoc local search optimization.
Notably, our results are competitive with classical heuristics like farthest insertion while offering an intrinsically different approach that leverages structural bias alone. This demonstrates that non-autoregressive, unsupervised methods can achieve effective performance on combinatorial optimization problems without the computational overhead of sequential decoding or iterative refinement. The detailed results are shown in Table 2.
@@ 8 Discussion and Future Work
In this paper, we use 24 configurations of $(\tau,\gamma)$ pairs for TSP-20 and TSP-50, and 8 configurations for TSP-100. This limited yet targeted hyperparameter exploration is sufficient to support our central claim: that structural inductive bias, when coupled with a permutation-based formulation, can drive the emergence of high-quality solutions in a fully unsupervised, non-autoregressive setting. We restrict our analysis to minimal hyperparameter settings and adopt the same architecture as UTSP Min et al. (2023), which is sufficiently expressive to illustrate our main claim. While preliminary evidence indicates that performance can be further improved through extensive hyperparameter tuning or architectural variations (e.g., alternative message passing schemes), such enhancements lie outside the scope of our primary contribution and are left for future work.
9 Appendix
Effectiveness of the Hamiltonian Cycle Ensemble
@@ 9 Appendix
10 Acknowledgement

### v3

@@ 6.1 Main Theorem
Consider the orbit of vertex 0 under this mapping:
@@
Sufficiency ($\Leftarrow$): Suppose $\gcd(k,n)=1$. Then by the argument above, the orbit of vertex 0 has size $\frac{n}{1}=n$. This means the sequence $\{0,k,2k,\ldots,(n-1)k\}$ modulo $n$ contains all distinct elements of $\{0,1,\ldots,n-1\}$.
@@
Figure 5: Best mean tour length on TSP instances of different sizes as the number of coprime shifts increases. Using more shift combinations significantly reduces tour length, outperforming the Greedy Multi-Start baseline. Both Christofides and Beam Search bounds are shown for comparison. From left to right: TSP-20, TSP-50, TSP-100.
@@ 6.2 Ensemble Training and Inference
Table 2: Detailed performance comparison of learning-based TSP solvers across different instance sizes (TSP-20/50/100). Metrics include average tour length and optimality gap (%). Results for baseline methods are taken from (Joshi et al., 2019). While all methods use uniformly generated TSP instances, test sets vary slightly across works. Note that many more recent models exist, we select a subset for comparison.
@@
PtrNet (Vinyals et al., 2015) SL, G 3.88 1.15% 7.66 34.48% - -
PtrNet (Bello et al., 2016) RL, G 3.89 1.42% 5.95 4.46% 8.30 6.90%
S2V (Khalil et al., 2017) RL, G 3.89 1.42% 5.99 5.16% 8.31 7.03%
GAT (Deudon et al., 2018) RL, G, 2-OPT 3.85 0.42% 5.85 2.77% 8.17 5.21%
GAT (Kool et al., 2018) RL, G 3.85 0.34% 5.80 1.76% 8.12 4.53%
GCN (Joshi et al., 2019) SL, G 3.86 0.60% 5.87 3.10% 8.41 8.38%
POMO (Kwon et al., 2020) RL,G 3.83 0.12% 5.73 0.64% 7.84 1.07%
@@
Hamiltonian cycle ensemble UL, NAR 3.97 3.52% 5.85 2.81% 8.14 5.03%
@@
Figure 5 demonstrates the effectiveness of this ensemble approach across 20, 50, and 100-node problems respectively. As the number of coprime shifts increases, the mean tour length decreases substantially, with dramatic improvements observed initially that gradually plateau. For TSP-20, using all $\varphi(20)=8$ coprime shifts reduces mean tour length from 4.06 to 3.97, surpassing both multi-start greedy and beam search performance. Similar trends are observed for TSP-50 and TSP-100, where the ensemble approach achieves mean tour lengths of 5.85 and 8.14 respectively when using all available coprime shifts. This demonstrates that leveraging multiple Hamiltonian cycle structures effectively mitigates the long tail problem while consistently improving solution quality. Here, each permutation learner in our framework is trained with respect to a fixed initial Hamiltonian cycle $\mathbb{V}^{k}$, which serves as a structural prior guiding solution formation. This initialization anchors the training process, focusing learning on the permutation of the initial Hamiltonian cycle $\mathbb{V}^{k}$, effectively biasing the model. Since different initial cycles encode distinct structural priors, using an ensemble of models with $\mathbb{V}^{k}$ promotes diversity and improves overall solution quality.
@@
Despite not using supervision or autoregressive decoding, our method achieves competitive results across all TSP sizes. Our Hamiltonian cycle ensemble approach significantly outperforms classical baselines, achieving optimality gaps of 3.52%, 2.81%, and 5.03% on TSP-20, TSP-50, and TSP-100 respectively, compared to greedy NN (all start cities) search’s 5.74%, 11.6%, and 14.8%. We also improve upon beam search variants as the problem size grows, with beam search achieving gaps of 3.13–3.92% on TSP-20, 17.0–17.9% on TSP-50, and 25.3–26.1% on TSP-100.
@@
Among learning-based methods, our approach demonstrates competitive performance. We achieve comparable results to Pointer Networks and S2V. Our method also performs competitively with supervised method (Joshi et al., 2019), achieving 5.03% optimality gap versus 8.38% on TSP-100. Our performance is comparable to the RL-based approaches. Notably, we are competitive with the GAT model of (Deudon et al., 2018) even when it is augmented with 2-OPT local search, a strong post-hoc refinement step. While models such as the attention-based approach by (Kool et al., 2018) leverage RL and autoregressive decoding, our unsupervised, non-autoregressive framework attains similar optimality gaps without requiring either RL training or explicit search procedures. However, we do not yet match the RL model such as (Kwon et al., 2020), which benefits from exploiting multiple equivalent solutions through parallel rollouts. Our results are also competitive with classical heuristics such as farthest insertion, while offering a fundamentally different approach grounded in structural inductive bias. Overall, our results highlight that non-autoregressive, unsupervised methods can effectively tackle combinatorial optimization problems without sequential decoding. Detailed comparisons are provided in Table 2.
In our experiments, we observe that it is not necessary to employ the full set of $\varphi(n)$ Hamiltonian cycles for effective ensembling. Instead, using a small subset can already yield strong approximations. For example, on TSP-100, selecting the shifts $k\in\{1,9,81,87,91\}$ achieves an average tour length of $8.18$, corresponding to an optimality gap of approximately $5.5\%$. Furthermore, the inference cost remains practical: whereas a single model requires about 0.400.40\,ms per instance, this five-model subset ensemble takes only 22\,ms in total, while still delivering significant improvements in robustness and solution quality.
@@ 9 Discussion and Future Work
In this paper, we use 24 configurations of $(\tau,\gamma)$ pairs for TSP-20, 6 configurations for TSP-50, and 3 configurations for TSP-100. This limited yet targeted hyperparameter exploration is sufficient to support our central claim: that structural inductive bias, when coupled with a permutation-based formulation, can drive the emergence of high-quality solutions in a fully unsupervised, non-autoregressive setting. We restrict our analysis to minimal hyperparameter settings and adopt the same architecture as UTSP (Min et al., 2023), which is sufficiently expressive to illustrate our main claim. While preliminary evidence indicates that performance can be further improved through extensive hyperparameter tuning or architectural variations (e.g., alternative message passing schemes), such enhancements lie outside the scope of our primary contribution and are left for future work.
Appendix A Effectiveness of the Hamiltonian Cycle Ensemble
@@
Appendix B Quadratic upper bound on the optimality gap
Theorem B.1 (Quadratic upper bound on the optimality gap).
Let $C(\mathbf{P}):=\langle\mathbf{D},\,\mathbf{P}\mathbb{V}\mathbf{P}^{\top}\rangle$ for a cost matrix $\mathbf{D}\in\mathbb{R}^{n\times n}$, a cyclic shift matrix $\mathbb{V}\in\mathbb{R}^{n\times n}$, and a permutation matrix $\mathbf{P}\in\Pi_{n}$. Let the set of optimal permutations be
$\mathcal{O}:=\arg\min_{\mathbf{P}\in\Pi_{n}}C(\mathbf{P}),\qquad C^{\star}:=\min_{\mathbf{P}\in\Pi_{n}}C(\mathbf{P}).$ (17)
Given a (soft) doubly-stochastic matrix $\mathbb{T}$ produced by the model and a hard permutation $\widehat{\mathbf{P}}$ obtained from $\mathbb{T}$ at inference, define
$\delta_{*}:=\min_{\mathbf{P}\in\mathcal{O}}\|\mathbb{T}-\mathbf{P}\|_{F},\qquad\varepsilon:=\|\widehat{\mathbf{P}}-\mathbb{T}\|_{F}.$ (18)
If $\|\mathbb{V}\|_{2}\leq 1$ and $\|\mathbb{T}\|_{2}\leq 1$, then
$C(\widehat{\mathbf{P}})-C^{\star}\ \leq\ \|\mathbf{D}\|_{F}\big(2\delta_{*}+\delta_{*}^{2}+2\varepsilon+\varepsilon^{2}\big).$ (19)
Proof.
Since $\Pi_{n}$ is finite, there exists $\mathbf{P}^{\dagger}\in\mathcal{O}$ such that $\delta_{*}=\|\mathbb{T}-\mathbf{P}^{\dagger}\|_{F}$. We decompose the gap into a “soft” part and a “rounding” part:
[…]
Soft term.
Let $E:=\mathbb{T}-\mathbf{P}^{\dagger}$. Expanding,
$\mathbb{T}\mathbb{V}\mathbb{T}^{\top}-\mathbf{P}^{\dagger}\mathbb{V}\mathbf{P}^{\dagger\top}=E\mathbb{V}\mathbf{P}^{\dagger\top}+\mathbf{P}^{\dagger}\mathbb{V}E^{\top}+E\mathbb{V}E^{\top}.$ (21)
Using the submultiplicative bounds
$\|AXB\|_{F}\leq\|A\|_{F}\|X\|_{2}\|B\|_{2},\qquad\|AXB\|_{F}\leq\|A\|_{2}\|X\|_{F}\|B\|_{2},$ (22)
together with $\|\mathbf{P}^{\dagger}\|_{2}=1$, $\|\mathbb{V}\|_{2}\leq 1$, and $\|E\|_{2}\leq\|E\|_{F}$, we obtain
$\|E\mathbb{V}\mathbf{P}^{\dagger\top}\|_{F}\leq\delta_{*},\quad\|\mathbf{P}^{\dagger}\mathbb{V}E^{\top}\|_{F}\leq\delta_{*},\quad\|E\mathbb{V}E^{\top}\|_{F}\leq\delta_{*}^{2}.$ (23)
Thus
$\|\mathbb{T}\mathbb{V}\mathbb{T}^{\top}-\mathbf{P}^{\dagger}\mathbb{V}\mathbf{P}^{\dagger\top}\|_{F}\leq 2\delta_{*}+\delta_{*}^{2}.$ (24)
By Cauchy–Schwarz,
$|C(\mathbb{T})-C(\mathbf{P}^{\dagger})|\ \leq\ \|\mathbf{D}\|_{F}(2\delta_{*}+\delta_{*}^{2}).$ (25)
Rounding term.
Let $\Delta:=\widehat{\mathbf{P}}-\mathbb{T}$, so $\|\Delta\|_{F}=\varepsilon$. Expanding,
$\widehat{\mathbf{P}}\mathbb{V}\widehat{\mathbf{P}}^{\top}-\mathbb{T}\mathbb{V}\mathbb{T}^{\top}=\Delta\mathbb{V}\mathbb{T}^{\top}+\mathbb{T}\mathbb{V}\Delta^{\top}+\Delta\mathbb{V}\Delta^{\top}.$ (26)
Using the same bounds and $\|\mathbb{T}\|_{2}\leq 1$, $\|\mathbb{V}\|_{2}\leq 1$,
$\|\Delta\mathbb{V}\mathbb{T}^{\top}\|_{F}\leq\varepsilon,\quad\|\mathbb{T}\mathbb{V}\Delta^{\top}\|_{F}\leq\varepsilon,\quad\|\Delta\mathbb{V}\Delta^{\top}\|_{F}\leq\varepsilon^{2},$ (27)
hence
$\|\widehat{\mathbf{P}}\mathbb{V}\widehat{\mathbf{P}}^{\top}-\mathbb{T}\mathbb{V}\mathbb{T}^{\top}\|_{F}\leq 2\varepsilon+\varepsilon^{2},$ (28)
and by Cauchy–Schwarz,
$|C(\widehat{\mathbf{P}})-C(\mathbb{T})|\ \leq\ \|\mathbf{D}\|_{F}(2\varepsilon+\varepsilon^{2}).$ (29)
Combine.
By the triangle inequality,
$C(\widehat{\mathbf{P}})-C^{\star}\ \leq\ \|\mathbf{D}\|_{F}\big(2\delta_{*}+\delta_{*}^{2}+2\varepsilon+\varepsilon^{2}\big).$ (30)
Lemma B.2 (Spectral norm of $\mathbb{V}$).
Let $\mathbb{V}\in\{0,1\}^{n\times n}$ be the cyclic shift matrix defined in Equation 3. Then
$\|\mathbb{V}\|_{2}=1.$ (31)
Proof.
The matrix $\mathbb{V}$ is a permutation matrix corresponding to a cyclic shift. Permutation matrices are orthogonal, i.e. $\mathbb{V}^{\top}\mathbb{V}=I$. Hence, all eigenvalues of $V$ have absolute value $1$, and
$\|\mathbb{V}\|_{2}=\sqrt{\lambda_{\max}(\mathbb{V}^{\top}\mathbb{V})}=\sqrt{\lambda_{\max}(I)}=1.$ (32)
Equivalently, $\mathbb{V}$ is diagonalizable by the discrete Fourier transform, with eigenvalues $\{e^{2\pi ik/n}:k=0,\dots,n-1\}$, all lying on the unit circle. Thus the spectral norm of $\mathbb{V}$ is exactly $1$. ∎
Lemma B.3 (Spectral norm of $\mathbb{T}$).
If $\mathbb{T}$ is doubly-stochastic, then $\|\mathbb{T}\|_{2}\leq 1$.
Proof.
By the Birkhoff–von Neumann theorem, any doubly-stochastic $\mathbb{T}$ can be written as a convex combination of permutation matrices: $\mathbb{T}=\sum_{k}\alpha_{k}\mathbf{P}_{k}$, with $\alpha_{k}\geq 0$ and $\sum_{k}\alpha_{k}=1$. The spectral norm is convex, hence
$\|\mathbb{T}\|_{2}=\Big\|\sum_{k}\alpha_{k}\mathbf{P}_{k}\Big\|_{2}\ \leq\ \sum_{k}\alpha_{k}\|\mathbf{P}_{k}\|_{2}\ =\ \sum_{k}\alpha_{k}\cdot 1\ =\ 1,$ (33)
since each permutation matrix $\mathbf{P}_{k}$ is orthogonal and thus has spectral norm $1$. ∎
Remark B.4 (Interpretation).
$\delta_{*}$ measures how close the learned soft matrix $\mathbb{T}$ is to some optimal permutation in $\mathcal{O}$, so the bound handles non-uniqueness naturally. When there are symmetries (e.g. reversed cycles, relabelings), $\delta_{*}$ will be the distance to the closest such symmetry, which tightens the bound compared to fixing an arbitrary $\mathbf{P}^{\dagger}$.
Appendix C Randomness by Hardware Perturbation Inference
We introduce Hardware Perturbation Inference (HPI), a simple yet effective technique that leverages the inherent non-determinism of low-level numerical operations to generate diverse inference outcomes without modifying the model or introducing explicit stochasticity. Even when using the same GPU architecture (e.g., NVIDIA H100), small numerical discrepancies can arise from differences in fused multiply–add (FMA) kernel execution and TensorFloat-32 (TF32) rounding modes. These subtle perturbations may propagate through the computation, leading to slightly different outputs. HPI exploits this phenomenon to produce multiple candidate solutions for the same problem instance, which can then be ensembled to improve robustness and solution quality—all without requiring changes to the model parameters or training procedure.
In our experiments, we apply HPI on NVIDIA H100 GPUs by toggling the use of FMA operations under TF32 precision. Specifically, we compare inference with TF32+FMA enabled versus disabled, which yields distinct perturbations in the numerical pathways and consequently different solutions for the same input instance $\mathbb{V}^{k}$. By combining these outputs in an ensemble, we observe further improvements in solution quality: on the TSP-100 benchmark, the ensemble reduces the optimality gap to 8.10.
While hardware-level perturbations provide a simple mechanism for generating diversity, there are many other ways to introduce randomness to further enhance ensemble performance. We leave a broader discussion of such strategies for future work.
Appendix D Zero-Shot Generalization
We propose a zero-shot evaluation strategy inspired by [Min and Gomes, 2025], leveraging dummy nodes. As an example, consider testing on a TSP instance with 95 cities using a model trained on TSP-100. To construct such a test case, we randomly select 5 parent nodes from the 95 cities and introduce 5 additional dummy nodes, each placed very close to one of the selected parents. This augmentation produces an effective 100-node instance, which we then solve using the TSP-100 model.
If the resulting tour connects each dummy node directly to its parent node, we merge them to recover a valid tour for the original 95-city problem. If this condition is not met, we repeat the process by re-sampling the parent and dummy nodes.
Using this strategy, our model achieves a mean tour length of $8.24$ across 1,0001{,}000 unseen test instances, compared to $9.49$ for the greedy baseline. For reference, the optimal mean tour length is $7.57$. These results demonstrate that our dummy-node construction enables effective zero-shot transfer across problem sizes while maintaining competitive performance.